# Iterating $\mathsf{PP}$-packages without Choice: A Cohen symmetric seed and a localization framework

### v3

@@ 5.2 Iteration-level diagonal automorphisms
Moreover, there is a group-theoretic splitting map $\mathrm{spl}:\mathscr{G}_{\alpha}\to\mathscr{G}_{\alpha+1}$ (the the splitting section) such that $\pi^{\alpha}_{\alpha+1}\circ\mathrm{spl}=\mathrm{id}_{\mathscr{G}_{\alpha}}$, and every $g\in\mathscr{G}_{\alpha+1}$ can be written as $g=\mathrm{spl}(h)\cdot k$ with $h=\pi^{\alpha}_{\alpha+1}(g)$ and $k\in K_{\alpha+1}$.
@@
  • At a successor stage $\alpha+1$, write $\mathbb{P}_{\alpha+1}\cong\mathbb{P}_{\alpha}*\dot{\mathbb{Q}}_{\alpha}$. In this paper, $\dot{\mathbb{Q}}_{\alpha}$ is (forced to be) a finite product of orbit packages
@@
When working at a fixed successor stage $\alpha+1$, we suppress the stage tag and write $\langle i,g\rangle$ for $\langle\alpha,i,g\rangle$ (so $\langle i,g\rangle\in\mathrm{Coords}_{\alpha+1}\setminus\mathrm{Coords}_{\alpha}$” means $i<n_{\alpha}$ and $g\in[f_{\alpha,i}]$).
We let $\operatorname{Sym}(\omega_{1})$ act on new coordinates by
@@
Fix a successor stage $\alpha+1$. Let $\pi\in\operatorname{Sym}(\omega_{1})$ and let $D^{\mathrm{new}}\subseteq\mathrm{Coords}_{\alpha+1}\setminus\mathrm{Coords}_{\alpha}$ be a set of package coordinates newly introduced at stage $\alpha+1$.
Assume $\pi$ fixes each $d\in D^{\mathrm{new}}$ under the induced action on tagged coordinates, i.e., if $d=\langle\alpha,i,g\rangle$ then $\pi\cdot d=d$ equivalently $\pi\cdot g=g$. (Here $\pi\cdot\langle\alpha,i,g\rangle:=\langle\alpha,i,\pi\cdot g\rangle$ using the orbit action on $f_{\alpha,i}]$.)
@@
Moreover, for fixed $D^{\mathrm{new}}$, the map $\pi\mapsto\kappa_{\alpha+1}(\pi,D^{\mathrm{new}})$ is a group homomorphism on its domain $\{\pi\in\operatorname{Sym}(\omega_{1}):\pi\cdot d=d\ \forall d\in D^{\mathrm{new}}\}$ (Here the action on new coordinates is $\pi\cdot\langle\alpha,i,g\rangle:=\langle\alpha,i,\pi\cdot g\rangle$.)
@@
If $(E,D)$ is globally admissible and $\pi\in\operatorname{Fix}(E)$, then $\widehat{\pi}^{\,D}\in\Delta_{\lambda}^{\uparrow}(E,D)$. Indeed, $\pi^{0}_{\lambda}(\widehat{\pi}^{\,D})=\pi\in\operatorname{Fix}(E)$ and $\widehat{\pi}^{\,D}$ fixes every coordinate in $D$ pointwise.
@@
This “characterization” definition makes monotonicity immediate (Lemma 5.34). In Lemma 5.44, we show that diagonal lifts $\widehat{\pi}^{\,D^{\prime}}$ (for $\pi\in\operatorname{Fix}(E)$ and $D^{\prime}\supseteq D$) lie in $\Delta_{\lambda}^{\uparrow}(E,D)$—this is immediate from the characterization. We do not need to prove that $\Delta_{\lambda}^{\uparrow}(E,D)$ is generated by such diagonal lifts; the characterization suffices for all filter-theoretic purposes.
@@
Let $g\in\mathscr{G}_{\lambda}$. Then:
@@
    The stage-0 projection $\pi^{0}_{\lambda}:\mathscr{G}_{\lambda}\to\operatorname{Sym}(\omega_{1})$ is a group homomorphism.
@@
    The element $g$ induces a bijection on the set of package coordinates at each stage $\beta<\lambda$. In particular, if $D$ is a countable set of package coordinates, then $g\cdot D$ is countable.
@@
    For $E\subseteq\omega_{1}$, define $g\cdot E:=\pi^{0}_{\lambda}(g)[E]$. Then $g\cdot E$ is countable whenever $E$ is countable.
@@
(i) This is part of the inverse/direct system structure of the iteration groups (cf. [5, Def. 3.1]) and the fact that $\pi^{0}_{\lambda}$ is defined stagewise and respects composition.
(ii) We prove by induction on stages that each $g\in\mathscr{G}_{\beta}$ determines a bijection of $\mathrm{Coords}_{\beta}$ (in the sense of Definition 5.22), and hence sends countable sets of coordinates to countable sets.
At stage 0 there are no package coordinates. At a successor stage $\beta+1$, write $g=\mathrm{st}(h)\cdot k$ as in Lemma 5.21, where $h=\pi^{\beta}_{\beta+1}(g)\in\mathscr{G}_{\beta}$ and $k\in K_{\beta+1}$ acts trivially on the $\mathbb{P}_{\beta}$-part. By the inductive hypothesis, $h$ induces a bijection of $\mathrm{Coords}_{\beta}$. The standard lift $\mathrm{st}(h)$ extends this action and fixes every element of $\mathrm{NewCoords}_{\beta+1}$. On the other hand, $k$ fixes $\mathrm{Coords}_{\beta}$ pointwise and acts on the successor forcing $\dot{\mathbb{Q}}_{\beta}$ by compositions of package automorphisms (Lemma 5.8) and diagonal-cancellation automorphisms (Definition 5.11), each of which induces a bijection on the new coordinate index set (and in particular cannot collapse distinct coordinates, by the admissibility hypothesis built into their definition). Therefore $g$ induces a bijection on $\mathrm{Coords}_{\beta+1}=\mathrm{Coords}_{\beta}\dot{\cup}\mathrm{NewCoords}_{\beta+1}$.
At a limit stage $\eta$, elements of $\mathscr{G}_{\eta}$ are (by definition) coherent direct-limit elements whose restrictions to earlier stages determine them uniquely (cf. [5, Def. 3.1]). The induced bijection on $\mathrm{Coords}_{\eta}=\bigcup_{\beta<\eta}\mathrm{Coords}_{\beta}$ is obtained by taking the union of the stagewise bijections, and is again a bijection.
Consequently, if $D$ is countable then so is its image under the induced bijection.
(iii) Immediate: $\pi^{0}_{\lambda}(g)\in\operatorname{Sym}(\omega_{1})$ is a bijection and hence preserves countability. ∎
@@
The group homomorphism property $\widehat{\pi}^{\,D}\circ\widehat{\sigma}^{\,D}=\widehat{\pi\sigma}^{\,D}$ follows coordinatewise: on $D$ both sides are the identity; on the Cohen base and on coordinates outside $D$, both sides agree with the standard action of $\pi\sigma$. ∎
Corollary 5.32 (Diagonal lifts swap Cohen reals).
Let $\widehat{\pi}^{\,D}\in\mathscr{G}_{\lambda}$ be a diagonal lift with $\pi^{0}_{\lambda}(\widehat{\pi}^{\,D})=\pi$. Then for each $\alpha<\omega_{1}$:
$\widehat{\pi}^{\,D}\cdot\dot{c}_{\alpha}=\dot{c}_{\pi(\alpha)},$
@@
In particular, if $\pi=(\beta\ \gamma)$ is a transposition with $\beta\neq\gamma$, then $\widehat{\pi}^{\,D}$ swaps $c_{\beta}\leftrightarrow c_{\gamma}$ in the generic extension.
@@
View $\dot{c}_{\alpha}$ as a $\mathbb{P}_{\lambda}$-name via the canonical inclusion of $\mathbb{P}_{0}$-names into $\mathbb{P}_{\lambda}$-names.
By Lemma 5.31(i), the diagonal lift $\widehat{\pi}^{\,D}$ restricts to $\pi$ on the Cohen base $\mathbb{P}_{0}=\operatorname{Add}(\omega,\omega_{1})$. The canonical name $\dot{c}_{\alpha}$ is defined by $\dot{c}_{\alpha}=\{(\check{n},p):p(\alpha,n)=1\}$. The action of $\pi$ on $\mathbb{P}_{0}$ is $(\pi\cdot p)(\pi(\alpha),n)=p(\alpha,n)$, which induces $\pi\cdot\dot{c}_{\alpha}=\dot{c}_{\pi(\alpha)}$. Since $\widehat{\pi}^{\,D}$ restricts to $\pi$ on stage 0, we have $\widehat{\pi}^{\,D}\cdot\dot{c}_{\alpha}=\dot{c}_{\pi(\alpha)}$. ∎
Remark 5.33 (Standard lifts as empty diagonal lifts).
The standard lift $\widehat{\pi}^{\,\emptyset}$ (Definition 5.27 with $D=\emptyset$) acts as $\pi$ on all coordinates, recovering the usual lifted action of $\operatorname{Sym}(\omega_{1})$ on the iteration. Thus the pushforward subgroups $\operatorname{Fix}(E)\leq\mathscr{G}_{\lambda}$ (acting via standard lifts) coincide with $\Delta_{\lambda}(E,\emptyset)$, and consequently $\Delta_{\lambda}^{\uparrow}(E,\emptyset)=\operatorname{Fix}(E)^{\uparrow}$.
This identification is crucial: all filter generators at limit stages—both Cohen-type and package-type—are of the form $\Delta_{\lambda}^{\uparrow}(E_{i},D_{i})$, with $D_{i}=\emptyset$ for Cohen generators and $D_{i}=\{f_{i}\}$ for package generators. The monotonicity lemma (Lemma 5.34) therefore applies uniformly to combine them.
@@
Let $g\in\Delta_{\lambda}^{\uparrow}(E_{1},D_{1})$. By the characterization definition:
  • $\pi^{0}_{\lambda}(g)\in\operatorname{Fix}(E_{1})\subseteq\operatorname{Fix}(E_{0})$ (since $E_{0}\subseteq E_{1}$ implies $\operatorname{Fix}(E_{1})\leq\operatorname{Fix}(E_{0})$).
  • $g$ fixes every coordinate in $D_{1}$ pointwise, hence fixes every coordinate in $D_{0}\subseteq D_{1}$ pointwise.
Thus $g\in\Delta_{\lambda}^{\uparrow}(E_{0},D_{0})$. ∎
Corollary 5.35 (Intersection of $\Delta^{\uparrow}$-groups).
For any family $\{(E_{i},D_{i}):i\in I\}$ (with $I$ finite or countable):
@@
Let $(E,D)$ be globally admissible at stage $\lambda$. For $\sigma\in\operatorname{Sym}(\omega_{1})$:
$\widehat{\sigma}^{\,\emptyset}\cdot\Delta_{\lambda}^{\uparrow}(E,D)\cdot(\widehat{\sigma}^{\,\emptyset})^{-1}=\Delta_{\lambda}^{\uparrow}(\sigma\cdot E,\sigma\cdot D).$
In particular, conjugates of diagonal generators remain diagonal generators (with shifted parameters).
@@
The proof is identical to Lemma 5.18, now applied at the iteration level. ∎
@@
Let $\lambda$ be a stage and $(E,D)$ be globally admissible at stage $\lambda$. For any $g\in\mathcal{G}_{\lambda}$, there is a countable $E^{\ast}\supseteq g\cdot E$ such that $(E^{\ast},\,g\cdot D)$ is globally admissible. Consequently $\Delta_{\lambda}^{\uparrow}(E^{\ast},\,g\cdot D)\subseteq g\,\Delta_{\lambda}^{\uparrow}(E,D)\,g^{-1}$.
@@
For any $g\in\mathscr{G}_{\lambda}$, any $E\subseteq\omega_{1}$, and any set $D$ of package coordinates:
$g\cdot\Delta_{\lambda}^{\uparrow}(E,D)\cdot g^{-1}=\Delta_{\lambda}^{\uparrow}(g\cdot E,g\cdot D)$
where $g\cdot E:=\pi^{0}_{\lambda}(g)[E]$ and $g\cdot D:=\{g\cdot d:d\in D\}$ is the image of $D$ under the action of $g$ on package coordinates.
@@
Fix $k\in\mathscr{G}_{\lambda}$ and write $\sigma:=\pi^{0}_{\lambda}(k)\in\operatorname{Sym}(\omega_{1})$. Let $g\in\Delta^{\uparrow}_{\lambda}(E,D)$, and put $\tau:=\pi^{0}_{\lambda}(g)\in\operatorname{Fix}(E)$.
Stage-0 part. Since $\pi^{0}_{\lambda}:\mathscr{G}_{\lambda}\to\operatorname{Sym}(\omega_{1})$ is a homomorphism, we have
$\pi^{0}_{\lambda}(kgk^{-1})=\sigma\tau\sigma^{-1}\in\operatorname{Fix}(\sigma[E])=\operatorname{Fix}(\pi^{0}_{\lambda}(k)[E]).$
Coordinate part. Let $d\in D$. By definition of $\Delta^{\uparrow}_{\lambda}(E,D)$, $g$ fixes $d$ pointwise. Using the coordinate action $\cdot$ of $\mathscr{G}_{\lambda}$ on $\mathrm{Coords}_{<\lambda}$ (Lemma 5.30), we compute
$(kgk^{-1})\cdot(k\cdot d)=k\cdot\bigl(g\cdot(k^{-1}\cdot(k\cdot d))\bigr)=k\cdot(g\cdot d)=k\cdot d,$
so $kgk^{-1}$ fixes every coordinate in $k\cdot D$ pointwise.
Thus $kgk^{-1}\in\Delta^{\uparrow}_{\lambda}(\pi^{0}_{\lambda}(k)[E],k\cdot D)$. The reverse inclusion follows by conjugating with $k^{-1}$, so equality holds. ∎
@@
Let $\beta<\lambda$ and let $D\subseteq\mathrm{Coords}_{\beta}$ (i.e., $D$ only uses coordinates introduced below stage $\beta$). Then:
@@
Let $g\in\Delta_{\lambda}^{\uparrow}(E,D)$. Then:
  • $\pi^{0}_{\lambda}(g)\in\operatorname{Fix}(E)$, and since $\pi^{0}_{\beta}\circ\pi^{\beta}_{\lambda}=\pi^{0}_{\lambda}$, we have $\pi^{0}_{\beta}(\pi^{\beta}_{\lambda}(g))=\pi^{0}_{\lambda}(g)\in\operatorname{Fix}(E)$.
  • $g$ fixes every coordinate in $D$ pointwise. Since $D\subseteq\mathrm{Coords}_{\beta}$, these coordinates are visible at stage $\beta$, and the projection $\pi^{\beta}_{\lambda}(g)$ also fixes them pointwise.
Hence $\pi^{\beta}_{\lambda}(g)\in\Delta_{\beta}^{\uparrow}(E,D)$. ∎

### v4

@@ 5.2 Iteration-level diagonal automorphisms
Moreover, there is a group-theoretic splitting map $\mathrm{spl}:\mathscr{G}_{\alpha}\to\mathscr{G}_{\alpha+1}$ (the splitting section) such that $\pi^{\alpha}_{\alpha+1}\circ\mathrm{spl}=\mathrm{id}_{\mathscr{G}_{\alpha}}$, and every $g\in\mathscr{G}_{\alpha+1}$ can be written as $g=\mathrm{spl}(h)\cdot k$ with $h=\pi^{\alpha}_{\alpha+1}(g)$ and $k\in K_{\alpha+1}$.
@@
Remark 5.22 (Projection maps are coherent).
For $\gamma<\beta$ we write $\pi^{\gamma}_{\beta}:\mathscr{G}_{\beta}\to\mathscr{G}_{\gamma}$ for the canonical projection homomorphisms of the iteration groups (cf. [5, Def. 3.1]). These satisfy coherence: for all $\gamma<\beta<\lambda$,
$\pi^{\gamma}_{\beta}\circ\pi^{\beta}_{\lambda}=\pi^{\gamma}_{\lambda}.$
Definition 5.23 (Stage-0 projection).
Set $\mathscr{G}_{0}:=\mathscr{G}=\operatorname{Sym}(\omega_{1})$. For each stage $\alpha$ define $\pi^{0}_{\alpha}:\mathscr{G}_{\alpha}\to\mathscr{G}$ by transfinite recursion:
$\pi^{0}_{0}=\operatorname{id},\qquad\pi^{0}_{\alpha+1}=\pi^{0}_{\alpha}\circ\pi^{\alpha}_{\alpha+1},$
and for limit $\lambda$ set
$\pi^{0}_{\lambda}(g):=\pi^{0}_{\alpha}\bigl(\pi^{\alpha}_{\lambda}(g)\bigr)\quad\text{for any }\alpha<\lambda.$
This is well-defined by Remark 5.22. In particular, for $\beta<\lambda$, $\pi^{0}_{\beta}\circ\pi^{\beta}_{\lambda}=\pi^{0}_{\lambda}$.
@@
  • At a successor stage $\alpha+1$, write $\mathbb{P}_{\alpha+1}\cong\mathbb{P}_{\alpha}*\dot{\mathbb{Q}}_{\alpha}$, where $\dot{\mathbb{Q}}_{\alpha}$ denotes the stage iterand (a finite product of orbit packages) and should not be confused with the localized $\mathsf{PP}$-package forcing $\mathbb{Q}_{f}$. In this paper, $\dot{\mathbb{Q}}_{\alpha}$ is (forced to be) a finite product of orbit packages
@@
When working at a fixed successor stage $\alpha+1$, we suppress the stage tag and write $\langle i,g\rangle$ for $\langle\alpha,i,g\rangle$ (so $\langle i,g\rangle\in\mathrm{Coords}_{\alpha+1}\setminus\mathrm{Coords}_{\alpha}$” means $i<m_{\alpha}$ and $g\in[f_{\alpha,i}]$).
We let $\mathscr{G}=\operatorname{Sym}(\omega_{1})$ act on new coordinates by
@@
Fix a successor stage $\alpha+1$. Let $\pi\in\mathscr{G}=\operatorname{Sym}(\omega_{1})$ and let $D^{\mathrm{new}}\subseteq\mathrm{Coords}_{\alpha+1}\setminus\mathrm{Coords}_{\alpha}$ be a set of package coordinates newly introduced at stage $\alpha+1$.
Assume $\pi$ fixes each $d\in D^{\mathrm{new}}$ under the induced action on tagged coordinates, i.e., if $d=\langle\alpha,i,g\rangle$ then $\pi\cdot d=d$, equivalently $\pi\cdot g=g$. (Here $\pi\cdot\langle\alpha,i,g\rangle:=\langle\alpha,i,\pi\cdot g\rangle$ using the orbit action on $f_{\alpha,i}]$.)
@@
Moreover, for fixed $D^{\mathrm{new}}$, the map
$\pi\longmapsto\kappa_{\alpha+1}(\pi,D^{\mathrm{new}})$
is a group homomorphism on its domain
$\{\pi\in\mathscr{G}=\operatorname{Sym}(\omega_{1}):\forall d\in D^{\mathrm{new}}\ (\pi\cdot d=d)\},$
where the action on new coordinates is the one fixed in Definition 5.24.
@@
Let $(E,D)$ be globally admissible at stage $\lambda$ and let $\pi\in\operatorname{Fix}(E)$. The diagonal lift $\widehat{\pi}^{\,D}$ is defined as an element of $\operatorname{Aut}(\mathbb{P}_{\lambda})$. Once we verify that $\widehat{\pi}^{\,D}\in\mathscr{G}_{\lambda}$ (Lemma 5.45), it follows immediately from the defining clause (A) that $\widehat{\pi}^{\,D}\in\Delta_{\lambda}^{\uparrow}(E,D)$.
@@
This “characterization” definition makes monotonicity immediate (Lemma 5.36). In Lemma 5.46, we show that once diagonal lifts $\widehat{\pi}^{\,D^{\prime}}$ are known to belong to $\mathscr{G}_{\lambda}$ (Lemma 5.45), then for $\pi\in\operatorname{Fix}(E)$ and $D^{\prime}\supseteq D$ we have $\widehat{\pi}^{\,D^{\prime}}\in\Delta_{\lambda}^{\uparrow}(E,D)$; this is then immediate from the characterization. We do not need to prove that $\Delta_{\lambda}^{\uparrow}(E,D)$ is generated by such diagonal lifts; the characterization suffices for all filter-theoretic purposes.
@@
Let $k\in\mathscr{G}_{\lambda}$. Then:
@@
    The stage-0 projection $\pi^{0}_{\lambda}:\mathscr{G}_{\lambda}\to\mathscr{G}$ is a group homomorphism.
@@
    The element $k$ induces a bijection on the set of package coordinates at each stage $\beta<\lambda$ (in the sense of Definition 5.24). In particular, if $D$ is a countable set of package coordinates, then $k\cdot D$ is countable.
@@
    For $E\subseteq\omega_{1}$, define $k\cdot E:=\pi^{0}_{\lambda}(k)[E]$. Then $k\cdot E$ is countable whenever $E$ is countable.
@@
(i) This is immediate from Definition 5.23 together with the coherence of the projections (Remark 5.22); cf. also [5, Def. 3.1].
(ii) We prove by induction on stages that each $k\in\mathscr{G}_{\beta}$ determines a bijection of $\mathrm{Coords}_{\beta}$, and hence sends countable sets of coordinates to countable sets.
At stage 0 there are no package coordinates. At a successor stage $\beta+1$, write $k=\mathrm{st}(h)\cdot u$ as in Lemma 5.21, where $h=\pi^{\beta}_{\beta+1}(k)\in\mathscr{G}_{\beta}$ and $u\in K_{\beta+1}$ acts trivially on the $\mathbb{P}_{\beta}$-part. By the inductive hypothesis, $h$ induces a bijection of $\mathrm{Coords}_{\beta}$. The standard lift $\mathrm{st}(h)$ extends this action and fixes every element of $\mathrm{NewCoords}_{\beta+1}$. On the other hand, $u$ fixes $\mathrm{Coords}_{\beta}$ pointwise and acts on the successor iterand $\dot{\mathbb{Q}}_{\beta}$ by compositions of package automorphisms (Lemma 5.8) and diagonal-cancellation automorphisms (Definition 5.11), each of which induces a bijection on the new coordinate index set. Therefore $k$ induces a bijection on $\mathrm{Coords}_{\beta+1}=\mathrm{Coords}_{\beta}\dot{\cup}\mathrm{NewCoords}_{\beta+1}$.
At a limit stage $\eta$, elements of $\mathscr{G}_{\eta}$ are coherent direct-limit elements whose restrictions to earlier stages determine them uniquely (cf. [5, Def. 3.1]). The induced bijection on $\mathrm{Coords}_{\eta}=\bigcup_{\beta<\eta}\mathrm{Coords}_{\beta}$ is obtained by taking the union of the stagewise bijections, and is again a bijection. Consequently, if $D$ is countable then so is $k\cdot D$.
(iii) Immediate, since $\pi^{0}_{\lambda}(k)\in\mathscr{G}=\operatorname{Sym}(\omega_{1})$ is a bijection and hence preserves countability. ∎
@@
The group homomorphism property $\widehat{\pi}^{\,D}\circ\widehat{\sigma}^{\,D}=\widehat{\pi\sigma}^{\,D}$ follows coordinatewise: on $D$ both sides are the identity; on the Cohen base and on coordinates outside $D$, both sides agree with the standard action of $\pi\sigma$.
Injectivity is immediate from (i): if $\widehat{\pi}^{\,D}=\widehat{\sigma}^{\,D}$ then their restrictions to the Cohen base agree, hence $\pi=\sigma$. ∎
Corollary 5.34 (Stage-0 projection controls the Cohen reals).
Let $g\in\mathscr{G}_{\lambda}$ and put $\pi:=\pi^{0}_{\lambda}(g)\in\mathscr{G}=\operatorname{Sym}(\omega_{1})$. Then for each $\alpha<\omega_{1}$,
$g\cdot\dot{c}_{\alpha}=\dot{c}_{\pi(\alpha)},$
@@
In particular, if $\pi=(\beta\ \gamma)$ is a transposition with $\beta\neq\gamma$, then $g$ swaps $c_{\beta}\leftrightarrow c_{\gamma}$ in the generic extension.
@@
View $\dot{c}_{\alpha}$ as a $\mathbb{P}_{\lambda}$-name via the canonical inclusion of $\mathbb{P}_{0}$-names into $\mathbb{P}_{\lambda}$-names. Recall
$\dot{c}_{\alpha}:=\{\langle\check{n},p\rangle:p(\alpha,n)=1\}.$
Since $\pi^{0}_{\lambda}(g)=\pi$, the restriction of $g$ to the Cohen base $\mathbb{P}_{0}=\operatorname{Add}(\omega,\omega_{1})$ is the usual $\pi$-action. Therefore $g\cdot\dot{c}_{\alpha}=\pi\cdot\dot{c}_{\alpha}=\dot{c}_{\pi(\alpha)}$. ∎
Remark 5.35 (Standard lifts and Cohen-type diagonal groups).
The standard lift $\widehat{\pi}^{\,\emptyset}$ from Definition 5.29(B) is the diagonal lift with no protected coordinates, hence it acts as $\pi$ on the Cohen base and as the standard $\pi$-action on every package coordinate.
For a countable $E\subseteq\omega_{1}$, the subgroup of standard lifts
$\operatorname{Fix}(E)^{\uparrow}:=\{\widehat{\pi}^{\,\emptyset}:\pi\in\operatorname{Fix}(E)\}\leq\mathscr{G}_{\lambda}$
is contained in the characterized diagonal group
$\Delta_{\lambda}^{\uparrow}(E,\emptyset)=\{g\in\mathscr{G}_{\lambda}:\pi^{0}_{\lambda}(g)\in\operatorname{Fix}(E)\},$
and in general this inclusion may be proper.
In the limit-stage diagonal filters, both Cohen-type constraints and package-coordinate constraints are handled uniformly using groups of the form $\Delta_{\lambda}^{\uparrow}(E,D)$ for countable parameters $E$ and countable sets of coordinates $D\subseteq\mathrm{Coords}_{<\lambda}$.
@@
Let $k\in\Delta_{\lambda}^{\uparrow}(E_{1},D_{1})$. By Definition 5.29(A):
  • $\pi^{0}_{\lambda}(k)\in\operatorname{Fix}(E_{1})\subseteq\operatorname{Fix}(E_{0})$ (since $E_{0}\subseteq E_{1}$ implies $\operatorname{Fix}(E_{1})\leq\operatorname{Fix}(E_{0})$);
  • $k$ fixes every coordinate in $D_{1}$ pointwise, hence fixes every coordinate in $D_{0}\subseteq D_{1}$ pointwise.
Thus $k\in\Delta_{\lambda}^{\uparrow}(E_{0},D_{0})$. ∎
Corollary 5.37 (Intersections contain a uniform diagonal subgroup).
Let $\{(E_{i},D_{i}):i\in I\}$ be a family with $I$ finite or countable, and assume each $E_{i}$ and each $D_{i}$ is countable. Then
@@
In particular, under $M\models\mathsf{DC}$ the unions on the right are countable, so the right-hand group is again of the form used to generate the diagonal stage filters.
@@
Let $E\subseteq\omega_{1}$ and let $D\subseteq\mathrm{Coords}_{<\lambda}$. For $\sigma\in\mathscr{G}=\operatorname{Sym}(\omega_{1})$,
$\widehat{\sigma}^{\,\emptyset}\cdot\Delta_{\lambda}^{\uparrow}(E,D)\cdot(\widehat{\sigma}^{\,\emptyset})^{-1}=\Delta_{\lambda}^{\uparrow}\bigl(\sigma[E],\ \widehat{\sigma}^{\,\emptyset}\cdot D\bigr),$
where $\widehat{\sigma}^{\,\emptyset}\cdot D:=\{\widehat{\sigma}^{\,\emptyset}\cdot d:d\in D\}$.
@@
Let $k\in\Delta_{\lambda}^{\uparrow}(E,D)$ and set $h:=\widehat{\sigma}^{\,\emptyset}\,k\,(\widehat{\sigma}^{\,\emptyset})^{-1}$. By Lemma 5.32(i), $\pi^{0}_{\lambda}$ is a homomorphism, hence
$\pi^{0}_{\lambda}(h)=\sigma\,\pi^{0}_{\lambda}(k)\,\sigma^{-1}.$
Since $\pi^{0}_{\lambda}(k)\in\operatorname{Fix}(E)$, the conjugate fixes $\sigma[E]$ pointwise, so $\pi^{0}_{\lambda}(h)\in\operatorname{Fix}(\sigma[E])$.
Next, for $d\in D$ we have $k\cdot d=d$, hence $h\cdot(\widehat{\sigma}^{\,\emptyset}\cdot d)=\widehat{\sigma}^{\,\emptyset}\cdot d$. Therefore $h$ fixes every coordinate in $\widehat{\sigma}^{\,\emptyset}\cdot D$ pointwise. By Definition 5.29(A), $h\in\Delta_{\lambda}^{\uparrow}(\sigma[E],\widehat{\sigma}^{\,\emptyset}\cdot D)$. The reverse inclusion follows by conjugating with $(\widehat{\sigma}^{\,\emptyset})^{-1}$. ∎
@@
Let $\lambda$ be a stage and let $(E,D)$ be globally admissible at stage $\lambda$. For any $k\in\mathscr{G}_{\lambda}$, there is a countable $E^{\ast}\supseteq k\cdot E$ such that $(E^{\ast},\,k\cdot D)$ is globally admissible.
Consequently,
$\Delta_{\lambda}^{\uparrow}(E^{\ast},\,k\cdot D)\subseteq k\,\Delta_{\lambda}^{\uparrow}(E,D)\,k^{-1}.$
Indeed, by Lemma 5.40 we have $k\,\Delta_{\lambda}^{\uparrow}(E,D)\,k^{-1}=\Delta_{\lambda}^{\uparrow}(k\cdot E,k\cdot D)$, and since $k\cdot E\subseteq E^{\ast}$ monotonicity (Lemma 5.36) yields $\Delta_{\lambda}^{\uparrow}(E^{\ast},k\cdot D)\leq\Delta_{\lambda}^{\uparrow}(k\cdot E,k\cdot D)$.
@@
Let $k\in\mathscr{G}_{\lambda}$, let $E\subseteq\omega_{1}$, and let $D\subseteq\mathrm{Coords}_{<\lambda}$. Then
$k\,\Delta_{\lambda}^{\uparrow}(E,D)\,k^{-1}=\Delta_{\lambda}^{\uparrow}(k\cdot E,k\cdot D),$
where $k\cdot E:=\pi^{0}_{\lambda}(k)[E]$ and $k\cdot D:=\{k\cdot d:d\in D\}$ (using the $\mathscr{G}_{\lambda}$-action on coordinates from Lemma 5.32).
@@
Fix $k\in\mathscr{G}_{\lambda}$ and put $\sigma:=\pi^{0}_{\lambda}(k)\in\mathscr{G}=\operatorname{Sym}(\omega_{1})$. Let $h\in\Delta^{\uparrow}_{\lambda}(E,D)$, and put $\tau:=\pi^{0}_{\lambda}(h)\in\operatorname{Fix}(E)$.
Stage-0 part. By Lemma 5.32(i), $\pi^{0}_{\lambda}$ is a homomorphism, hence
$\pi^{0}_{\lambda}(khk^{-1})=\sigma\tau\sigma^{-1}\in\operatorname{Fix}(\sigma[E])=\operatorname{Fix}(\pi^{0}_{\lambda}(k)[E]).$
Coordinate part. Let $d\in D$. By definition of $\Delta^{\uparrow}_{\lambda}(E,D)$, $h$ fixes $d$ pointwise. Using the coordinate action of $\mathscr{G}_{\lambda}$ on $\mathrm{Coords}_{<\lambda}$ (Lemma 5.32),
$(khk^{-1})\cdot(k\cdot d)=k\cdot\bigl(h\cdot(k^{-1}\cdot(k\cdot d))\bigr)=k\cdot(h\cdot d)=k\cdot d,$
so $khk^{-1}$ fixes every coordinate in $k\cdot D$ pointwise.
Thus $khk^{-1}\in\Delta^{\uparrow}_{\lambda}(k\cdot E,k\cdot D)$, proving $k\,\Delta_{\lambda}^{\uparrow}(E,D)\,k^{-1}\subseteq\Delta^{\uparrow}_{\lambda}(k\cdot E,k\cdot D)$. The reverse inclusion follows by conjugating with $k^{-1}$. ∎
@@
Let $\beta<\lambda$, let $E\subseteq\omega_{1}$, and let $D\subseteq\mathrm{Coords}_{\beta}$ (i.e., $D$ only uses coordinates introduced below stage $\beta$). Then
@@
Let $k\in\Delta_{\lambda}^{\uparrow}(E,D)$. Then:
  • $\pi^{0}_{\lambda}(k)\in\operatorname{Fix}(E)$, and by Definition 5.23 (using coherence from Remark 5.22) we have $\pi^{0}_{\beta}\circ\pi^{\beta}_{\lambda}=\pi^{0}_{\lambda}$. Hence
    $\pi^{0}_{\beta}(\pi^{\beta}_{\lambda}(k))=\pi^{0}_{\lambda}(k)\in\operatorname{Fix}(E).$
  • $k$ fixes every coordinate in $D$ pointwise. Since $D\subseteq\mathrm{Coords}_{\beta}$, these coordinates are already present at stage $\beta$, and the projection $\pi^{\beta}_{\lambda}(k)$ acts on $\mathrm{Coords}_{\beta}$ as the restriction of $k$ (cf. Lemma 5.32(ii)), so $\pi^{\beta}_{\lambda}(k)$ fixes every coordinate in $D$ pointwise.
Hence $\pi^{\beta}_{\lambda}(k)\in\Delta_{\beta}^{\uparrow}(E,D)$ by Definition 5.29(A). ∎